OpenStudy (anonymous):

Give exact and approximate solutions to three decimal places (x-10)^2=12

OpenStudy (anonymous):

\[10-2\sqrt{3},10+2\sqrt{3}\] 6.536,13.464???

sam (.sam.):

\[(x-10)^2=12\] square root both sides \[x-10=\pm \sqrt{12}\] ---------------------------------- \[x-10= \sqrt{12}~~~~~~and~~~~~~~x-10= -\sqrt{12}\] Solve both equations

sam (.sam.):

yes, you're correct

OpenStudy (anonymous):

ty